# Epidemic Forecasting with a Hybrid Deep Learning Method Using CNN-LSTM With WOA-GWO Parameter Optimization: Global COVID-19 Case Study

### v2

@@ 2.3.1 Grey wolf optimization algorithm
The Grey Wolf Optimizer (GWO) is a metaheuristic optimization algorithm inspired by the hunting behavior of grey wolves. Proposed by [mirjalili2014grey], GWO mimics the social hierarchy within wolf packs to solve optimization problems. The GWO algorithm involves the following steps:
@@ 2.3.2 Whale optimization algorithm
The Whale Optimization Algorithm (WOA) draws inspiration from the Bubble-net hunting strategy of whales [mirjalili2016whale]. WOA comprises three main steps: encircling, exploitation, and exploration.
@@ 2.3.3 Proposed Hybrid Metaheuristic Optimization Algorithm
1 $Initialize$ population of $n$ agents $X_{i}\ (i=1,2,…,n)$ ;
2 $Initialize$ parameters $a,A,C,l,$ and $p$;
@@ 6 $X_{\delta}$ = third-best agent;
7 while $t\leq$ number of iterations do
@@ 3 Forecasting Framework
The proposed method combines the CNN and LSTM networks for forecasting the number of infected, cured, and dead people. As described in 2.1, the CNNs are able to extract local and deep features using their convolutional layers. On the other hand, the LSTM models can capture long-term dependencies, making them a good choice for sequence modeling. As a result, their combination can improve the accuracy of prediction [guo2020short].
@@ 4 Train LSTM model on $\mathbf{X}_{\text{cnn}}$ with hyperparameter tuning;
5 for $i$ in range(number of forecast steps) do
@@ 9       Append $\mathbf{y}_{\text{forecast}}$ to $\mathbf{X}_{\text{cnn}}$;
return Forecasted values from $\mathbf{X}_{\text{cnn}}$
@@ 5.1 Evaluation Setup and Metrics
The proposed methodology was implemented using Python version 3.8 on the Google Colab  [^1] https://colab.research.google.com/ platform, equipped with a K80 GPU and 12 GB of RAM. The implementation utilized the NiaPy [NiaPyJOSS2018] and Keras [chollet2015keras] libraries for incorporating the metaheuristic algorithms and deep learning models. To account for the stochastic nature of deep learning networks and their tendency to yield varying results with each run, the algorithms were executed ten times, and the average outcomes were subsequently reported. The batch size was set to 1, and the number of epochs was set to 100.
@@ 5.2 Hyperparameter Tuning
As indicated in Section 3.3, the proposed approach employs the GWO-WOA to optimize the hyperparameters of the CNN-LSTM model. These hyperparameters include the number of filters, kernel size of the convolutional layer, pool size of the Max-pooling layer, and the number of units in the LSTM layer. Furthermore, a comparative analysis was conducted to evaluate the performance of GWO-WOA in comparison to other optimization algorithms, namely Genetic Algorithm (GA), GWO [mirjalili2014grey], and WOA [mirjalili2016whale]. The configuration settings for these algorithms, as proposed in the NiaPy library, are presented in Table 1.
@@ 5.3 Comparison With Other Methods
- Able to learn local features - Limited ability to capture long-term dependencies [han2022survey]
@@
- Effective in modeling sequential data with spatial relationships - Requires more training data[rahman2019deep]
@@
- Less prone to overfitting - Less expressive than LSTM [shah2021deep]
@@
- Effective in modeling sequential data - Higher memory requirements[wozniak2023bilstm]
@@
- Effective in capturing spatial and temporal dependencies - May require more training data[lindemann2021survey]
@@ 5.4 Statistical Comparison of Different Methods
In this section, the Friedman test and the Nemenyi post-hoc test are used for more precise evaluation to distinguish the performance of different forecasting methods. This test is non-parametric and does not hypothesize normally distributed values for the sample[nemenyi1963distribution]. The null hypothesis for the conducted Friedman test concludes that all methods are equivalent and there is no significant difference between them by having similar rank. In the Nemenyi post-hoc test if the average ranking for two different methods is greater than the critical difference their performances are considerably different. The average ranks by the Friedman test are presented in Table 3.
@@
To calculate the Friedman statistic the following equation is used [nemenyi1963distribution]:
@@
$N$ is the number of tests carried out and k and $R_{i}$ denote the number of methods and their ranks, respectively. By considering the significance value of 0.05, the critical value for 6 different approaches in 24 different tests(countries) is approximately 3.146. Consequently, since the Friedman statistic of 5.5 is greater than the critical value of 3.146, the null hypothesis is rejected and all methods do not have equivalent performance. Therefore, the Nemenyi test can be conducted to compare the approaches as their performance concluded to have significant differences. In order to perform the Nemenyi post-hoc test, the following equation is considered:

### v3

@@ 2.3.1 Grey wolf optimization algorithm
The Grey Wolf Optimizer (GWO) is a metaheuristic optimization algorithm inspired by the hunting behavior of grey wolves. Proposed by [36], GWO mimics the social hierarchy within wolf packs to solve optimization problems. The GWO algorithm involves the following steps:
@@ 2.3.2 Whale optimization algorithm
The Whale Optimization Algorithm (WOA) draws inspiration from the Bubble-net hunting strategy of whales [35]. WOA comprises three main steps: encircling, exploitation, and exploration.
@@ 2.3.3 Proposed Hybrid Metaheuristic Optimization Algorithm
1
21exInitializeInitialize population of $n$ agents $X_{i}\ (i=1,2,…,n)$ ;
3 InitializeInitialize parameters $a,A,C,l,$ and $p$;
@@ 7 $X_{\delta}$ = third-best agent;
8
91exwhile $t\leq$ number of iterations do
@@ 3 Forecasting Framework
The proposed method combines the CNN and LSTM networks for forecasting the number of infected, cured, and dead people. As described in 2.1, the CNNs are able to extract local and deep features using their convolutional layers. On the other hand, the LSTM models can capture long-term dependencies, making them a good choice for sequence modeling. As a result, their combination can improve the accuracy of prediction [19].
@@ 5 Train LSTM model on $\mathbf{X}_{\text{cnn}}$ with hyperparameter tuning;
6
71exfor $i$ in range(number of forecast steps) do
@@ 11   Append $\mathbf{y}_{\text{forecast}}$ to $\mathbf{X}_{\text{cnn}}$;
14
1exreturn Forecasted values from $\mathbf{X}_{\text{cnn}}$
@@ 5.1 Evaluation Setup and Metrics
The proposed methodology was implemented using Python version 3.8 on the Google Colab  [^1] https://colab.research.google.com/ platform, equipped with a K80 GPU and 12 GB of RAM. The implementation utilized the NiaPy [61] and Keras [11] libraries for incorporating the metaheuristic algorithms and deep learning models. To account for the stochastic nature of deep learning networks and their tendency to yield varying results with each run, the algorithms were executed ten times, and the average outcomes were subsequently reported. The batch size was set to 1, and the number of epochs was set to 100.
@@ 5.2 Hyperparameter Tuning
As indicated in Section 3.3, the proposed approach employs the GWO-WOA to optimize the hyperparameters of the CNN-LSTM model. These hyperparameters include the number of filters, kernel size of the convolutional layer, pool size of the Max-pooling layer, and the number of units in the LSTM layer. Furthermore, a comparative analysis was conducted to evaluate the performance of GWO-WOA in comparison to other optimization algorithms, namely Genetic Algorithm (GA), GWO [36], and WOA [35]. The configuration settings for these algorithms, as proposed in the NiaPy library, are presented in Table 1.
@@ 5.3 Comparison With Other Methods
- Able to learn local features - Limited ability to capture long-term dependencies [21]
@@
- Effective in modeling sequential data with spatial relationships - Requires more training data[45]
@@
- Less prone to overfitting - Less expressive than LSTM [50]
@@
- Effective in modeling sequential data - Higher memory requirements[65]
@@
- Effective in capturing spatial and temporal dependencies - May require more training data[33]
@@ 5.4 Statistical Comparison of Different Methods
In this section, the Friedman test and the Nemenyi post-hoc test are used for more precise evaluation to distinguish the performance of different forecasting methods. This test is non-parametric and does not hypothesize normally distributed values for the sample[40]. The null hypothesis for the conducted Friedman test concludes that all methods are equivalent and there is no significant difference between them by having similar rank. In the Nemenyi post-hoc test if the average ranking for two different methods is greater than the critical difference their performances are considerably different. The average ranks by the Friedman test are presented in Table 3.
@@
To calculate the Friedman statistic the following equation is used [40]:
@@
NN is the number of tests carried out and k and $R_{i}$ denote the number of methods and their ranks, respectively. By considering the significance value of 0.05, the critical value for 6 different approaches in 24 different tests(countries) is approximately 3.146. Consequently, since the Friedman statistic of 5.5 is greater than the critical value of 3.146, the null hypothesis is rejected and all methods do not have equivalent performance. Therefore, the Nemenyi test can be conducted to compare the approaches as their performance concluded to have significant differences. In order to perform the Nemenyi post-hoc test, the following equation is considered: